Corbin Co. forecasts the following nominal cash flows for a project:

Year.......Cash Flows

0............-2,000

1............1,250

2............750

3...........500

The real rate of inflation is 7% and the inflation rate is 4%. Calculate the NPV for the project. Round to 2 decimals.
Answer #1

nominal rate=(1+real rate)*(1+inflation)-1=1.07*1.04-1=11.2800000%

NPV=-2000+1250/1.1128+750/1.1128^2+500/1.1128^3=91.79292